Solve each of the system by using either addition or substitution. If a unique solution does not exist, state whether the system is dependent or inconsistent.
10x + 2y = 7
y = -5x + 3
Any help is appreciated. Thanks

Answer by checkley71(8403)

10x+2y=7 & y=-5x+3 multiply equation #2 by -2 & add
-10x-2y=-6
-------------------------
0x+0y=1 inconsistant